# A note on thermal effects in non-linear models for plasma-based acceleration

### v1

@@ 2.3 Blowout radius equation
Figure 4: Radial cross section of the source term $S=-\rho-J_{z}$ for different temperature realizations, at a fixed value of the co-moving coordinate $\xi_{0}=7.0$, corresponding to the longitudinal center of the Gaussian bunch. As the background temperature increases, the electron accumulation region at the $r_{b}$ interface broadens and its peak amplitude decreases. This behavior highlights the need for a temperature dependent parametrization of the source term $S$ through a suitably defined $\Delta(\mu_{i})$.
We now have all the ingredients for the determination of an equation of motion defining the interface of the bubble. If Eq.˜15 is evaluated at $r=r_{b}$, the electromagnetic field expressions from Section˜2.2 can be used, and the only remaining task is to find an appropriate parametrization of $\psi_{0}$. This can be done by integrating Eq.˜18 after recurring again to the single-sheath parametrization of the source term used in [8, 9]:
@@
where $n_{\Delta}$ can be obtained from the continuity equation (see Eq.˜17). One obtains:
@@
In Lu’s model [8, 9] (that can be recovered by setting $E^{(0)}=1$, $p_{z}^{(0)}=0$ in Eqs.˜30 and 32), the evolution of the bubble radius $r_{b}(\xi)$ is governed by an ordinary differential equation (ODE), whose source term encodes the influence of the electron bunch. In the cold plasma regime, this equation is fully deterministic. However, when thermal effects are introduced, the model becomes stochastic: the ODE’s coefficients Eqs.˜30 and 32 become effectively random due to thermal perturbations in the plasma, and this in turn originates into random initial conditions for the ODE. As a result, the integration of the ODE must be interpreted over an ensemble of particles, and one has to shift from the deterministic view of the cold model, where the bubble radius $r_{b}(\xi)$ was representing the "trajectory of the innermost electron", to a warm interpretation where the bubble radius is now a statistical average $\langle r_{b}(\xi)\rangle$ of an ensemble of randomly sampled trajectories. This stochasticity is driven by the nature of thermal plasma electrons, that even at rest possess non-zero momenta that can be sampled from thermal equilibrium distributions.
@@ 3 Model Characterization and Comparison with Numerical Simulations
Figure 5: Parametric scan of the electron sheath width parameters $(\epsilon,\Delta_{l})$, shown through color maps of the relative error metric $C(\epsilon,\Delta_{l})$, as defined in Eq.˜37. Each panel corresponds to a different background temperature: $k_{b}T_{i}=0,1$ and $2\;\rm{keV}$ (panels from top to bottom). The metric quantifies the agreement between the bubble radius $r_{b}^{(sim)}$ extracted from PIC simulations and the theoretical prediction $r_{b}^{(th)}(\epsilon,\Delta_{l})$. Minima in the plots indicate the most suitable parametrization in order to reproduce the simulation results at each temperature.
@@
In this section, we will detail the strategy to solve numerically the blowout radius equation Eq.˜29, and all the considerations that have to be operated in the presence of a warm background plasma. Furthermore, we will present comparisons with PIC simulations performed using FBPIC [40], a quasi-3D solver that employs a pseudo-spectral decomposition of the fields in Fourier-Hankel space, enabling tunable azimuthal resolutions, and which also includes standard PIC components for particle evolution like particle pushers [41] and charge deposition schemes [42, 43, 44]. All numerical results presented consider a domain of size $L_{r}=3.0$ and $L_{z}=15.0$, with cell resolutions $\Delta_{z}=\Delta_{r}=0.01$, and computational time step $\Delta t=0.001$. The background plasma is modeled using $16$ particles per cell. We adopt the bi-Gaussian bunch described in Eq.˜1, with $\alpha=84$, $\sigma_{r}=0.1$, and $\sigma_{z}=\sqrt{2}$, in order to satisfy the narrow bunch condition described at the end of Section˜2.2. The choice of parameters grants a complete blowout regime (the normalized charge parameter  [^2] For the bi-Gaussian bunch detailed in Eq. 1, one defines the normalized charge parameter as $\tilde{Q}=\alpha(2\pi)^{3/2}\sigma_{r}^{2}\sigma_{z}$ [45]. , $\tilde{Q}$, which is often used to evaluate linearity regimes in PWFA applications, equals $\sim 18$) and adhere to the parameters chosen in [8, 9]. In Fig.˜1, we show a snapshot of the system under consideration coming from a PIC simulation. The upper half of the figure corresponds to a cold plasma setup, while the lower half shows the warm plasma case. A clear difference emerges between the two: in the presence of a finite plasma temperature, the plasma bubble becomes longitudinally smaller. This reduction in bubble size is due to thermal effects on the plasma wave frequency [33, 34]. This effect is characterized more quantitatively through the observables $\ell_{\parallel}$ and $\ell_{\bot}$ defined in Fig.˜1, which measure the spatial extent of the interface $r_{b}(\xi)$, the boundary separating the interior of the bubble from the surrounding plasma sheath, where electrons accumulation occurs. In warm plasmas, this sheath is broader and less sharply defined compared to the cold case. This is a direct consequence of the statistical spread in electron momenta due to thermal fluctuations.
@@
In order to investigate how the theoretical model exposed in Section˜2 captures this behavior, we numerically solve the ODE in Eq.˜29 for an ensemble of $500$ stochastic realizations of the initial quantities $E^{(0)}$ and $p_{z}^{(0)}$. These quantities are sampled according to a thermal distribution, i.e., a normal distribution $\mathcal{N}(\nu,\sigma^{2})$ with mean $\nu$ and variance $\sigma^{2}$, using the relations:
@@
This sampling reflects first on the form of the coefficients of the ODE Eqs.˜30 and 32. Furthermore, each sampled state yields a different initial condition for the trajectory solver. From Eq.˜12 one gets:
@@
using the same parameterization as in Lu’s model [8, 9], in order to evaluate temperature effects independently of possible changes due to a different choice for the source term $S$. In the panel (a) of Fig.˜2, we show the solution of Eq.˜29 for an initial plasma temperature of $k_{b}T_{i}=1\rm{keV}$. The stochasticity of the initial conditions gives rise to an ensemble of trajectories (shown in gray), from which we define the representative bubble trajectory as their statistical average, $\langle r_{b}(\xi)\rangle$ (shown in orange). In panel (b), we compare results coming from a cold realization of the model with the warm $k_{b}T_{i}=1\rm{keV}$ case. One observes that the statistical average $\langle r_{b}(\xi)\rangle$ (dashed orange line) shows only modest deviations from the cold bubble trajectory (shown in blue), indicating that the averaging process alone is not sufficient to fully capture thermal effects, at least at the temperature values considered here. Nevertheless, the stochastic fluctuations that arise from the ODE are not negligible. In panel (c) we quantify these fluctuations through
@@
evaluated at $\xi=11.0$, and computed for different values of the initial background temperature. As anticipated from Section˜3, the fluctuations exhibit a $\sim(k_{b}T_{i})^{1/2}$ scaling, reminiscent of the behavior found in the expression for the Deby length. We note however that this analogy is only formal, as Debye shielding appears only in the presence of thermal equilibration, and such effect does not appear in this context. A similar analysis is presented in Fig.˜3 for the first derivative $r_{b}^{\prime}$, which can be related via Eq.˜12 to momentum variables, and thus its ensemble fluctuations are naturally connected to thermal spread. Panels (a) and (b) lead to the same conclusions as in the previous figures: with the parametrization of Eq.˜35, the stochastic averages deviate only slightly from the cold solution. On the other hand the fluctuations remain significant at both the head and tail of the bubble, and when plotting the metric $\Delta r_{b}^{\prime}(\xi)$, defined analogously to Eq.˜36, we observe a trend that echoes the temperature profile reported in mono-dimensional studies [46], with a rise at the bubble front and an even more pronounced peak at its tail.
@@
The presence of thermal fluctuations suggests a non-trivial impact on the structure of the electron sheath that defines the edge of the bubble and hence, on the effective source term itself. Such effect is clearly visible in Fig.˜4, where we show results from PIC simulations detailing a radial cross section of the source term $S$ at a fixed value of the co-moving coordinate, $\xi_{0}=7.0$, corresponding to the longitudinal center of the bi-Gaussian bunch. As the background temperature increases, the region of electron accumulation at the boundary of the bubble broadens and its peak amplitude decreases. Indeed, this behavior suggests that an accurate description of the warm plasma regime requires not only the inclusion of stochastic initial conditions in the coefficients $E^{(0)}$ and $p_{z}^{(0)}$, but also in a temperature dependent parametrization of the source term that captures the underlying statistical spread induced by thermal effects.
Figure 6: Results of the numerical integration of the ODE in Eq.˜29 for two different temperature realizations, compared against the corresponding PIC simulations. In the cold case (top half of the figure), the trajectory $r_{b}(\xi)$ is deterministic, while in the warm case we plot the statistical average $\langle r_{b}(\xi)\rangle$ derived from an ensemble of $500$ stochastic realizations of the ODE. In both cases, the electron sheath widths $\Delta$ adopted in the resolution of the model are chosen following the parameter scan depicted in Fig.˜5.
A derivation of such parametrization from first principles is indeed challenging and worth of future investigations. As a preliminary step in this direction, we have performed a parameter-space scan over the terms $(\epsilon,\Delta_{l})$ defined in Eq.˜35 to identify, a posteriori, the most suitable representation of the electron sheath width $\Delta$, and hence of the source term $S$. To this end, we have introduced a metric that quantifies the discrepancy between the bubble trajectory $r_{b}^{(sim)}$ obtained from PIC simulations and a theoretical realization $\langle r_{b}\rangle(\epsilon,\Delta_{l})$ of the model Eq.˜29, corresponding to a given set of parameters:
@@
The results of this parametric scan are shown in Fig.˜5 for three values of background temperature, $k_{b}T_{i}=0,1,2\;\rm{keV}$. Comparisons of the $10\%$ error curves (depicted as dashed white lines in the figure) show that an increase in temperature implies a change in the optimal values of both $\Delta_{l}$ and $\epsilon$. The single-sheath parametrization of the source term adopted by [8, 9] and in this work implies that the contribution of the source term $S$ to the pseudo-potential $\psi$ can be effectively divided into three distinct regions: an inner cavity, a narrow electron sheath, and an outer linear-response region where plasma perturbation is minimal. One then sees that as $k_{b}T_{i}$ increases, the $10\%$ error curves marking the optimal parameters region shifts downward. This shows that the optimal $\Delta_{l}$, expected to be on the order of the cold plasma wavelength $k_{p}^{-1}$ (equal to $1$ in dimensionless coordinates [8, 9]), should slightly decrease. This trend can be interpreted as a consequence of the overall compression of characteristic length scales in warm plasmas compared to cold ones. Since $\Delta_{l}$ represents the characteristic length of the linear response region [8, 9], a slight thermal contraction is therefore expected. Conversely, the parameter $\Delta_{s}=\epsilon r_{b}$ sets the width of the sheath region, where the source term rises sharply. As shown in Fig.˜4, the height of this region diminishes with increasing temperature; charge conservation then requires a compensating increase in its width. In Fig.˜5, this is reflected into a "thermal shrinkage" of the $10\%$ error curves at low $\epsilon$, effectively excluding these values from the optimal region. Overall, as the background temperature increases, the contribution from the sheath region becomes progressively less significant (see Fig.˜4). Accordingly, a reduction in the $\Delta$ parameter as a whole is expected to reflect this diminished role.
Figure 7: Longitudinal ($\ell_{\parallel}$, top panel) and transverse ($\ell_{\bot}$, bottom panel) bubble lengths as functions of the background temperature. Results coming from PIC simulations (in orange) are compared to the ones coming from the model (in blue). The model captures the qualitative decrease in wavelength induced by thermal pressure, as well as a measurable contraction in the transverse size. The transversal bubble length $\ell_{\bot}$ has been captured at $\xi=11.5$, and results may vary depending on the selected value. The PIC measures have been obtained by averaging an ensemble of different simulations, in order to minimize numerical noise.
Fig.˜6 shows that the combined effect of the new source term parametrization and the stochastic averaging procedure introduced earlier yields a reasonable reconstruction of the interface $r_{b}$ in the warm-plasma case (bottom half of the figure). In particular, the model successfully captures the reduction in the longitudinal extent of the bubble relative to the cold-plasma case (top half), a consequence of thermal effects. This trend is more clearly illustrated in Fig.˜7, where the two previously defined observables $\ell_{\parallel}$ and $\ell_{\bot}$, corresponding to the longitudinal and transversal bubble lengths, are plotted as functions of the background temperature. The model reproduces the qualitative decrease in wavelength driven by thermal pressure and also reflects a measurable contraction in transverse size. Overall, the magnitude of these effects may increase for more linear regimes (corresponding to smaller values of $\tilde{Q}$). However, in such cases, the assumptions underlying the model presented in this work, namely the requirement of a complete blowout, would no longer hold. We therefore leave the assessment of this thermal model’s performance in a more linear wakefield regime to future studies. Lastly, in Fig.˜8 we report the values of the on-axis longitudinal electric fields, for both a cold (top panel) and a warm ($k_{b}T_{i}=1\;\rm{keV}$, bottom panel) background plasma. PIC results are confronted against solutions of the theoretical model presented in this article (that can easily be computed from $r_{b}^{\prime}$ [8, 9]). It is observed that the model is able to reproduce accurately the field inside the bubble, with noticeable deviations only at its tail (where the interface $r_{b}(\xi)$ becomes non derivable). Furthermore, the structure of the longitudinal field is not significantly altered by the warm background plasma, as already expected from previous studies [22].

### v2

@@ 2.3 Blowout radius equation
Figure 4: Radial cross section of the source term $S=-\rho-J_{z}$ for different temperature realizations, at a fixed value of the co-moving coordinate $\xi_{0}=7.0$, corresponding to the longitudinal center of the Gaussian bunch.
We now have all the ingredients for the determination of an equation of motion defining the interface of the bubble. If Eq.˜14 is evaluated at $r=r_{b}$, the electromagnetic field expressions from Section˜2.2 can be used, and the only remaining task is to find an appropriate parametrization of $\psi_{0}$. This can be done by integrating Maxwell’s equations after recurring again to the single-sheath parametrization of the source term used in [32, 30]:
@@
where $n_{\Delta}$ can be obtained from the charge continuity equation. One obtains:
@@
In Lu’s model [32, 30] (that can be recovered by setting $E^{(0)}=1$, $p_{z}^{(0)}=0$ in Eqs.˜21 and 23), the evolution of the bubble radius $r_{b}(\xi)$ is governed by an ordinary differential equation (ODE), whose source term encodes the influence of the electron bunch. In the cold plasma regime, this equation is fully deterministic. However, when thermal effects are introduced, the model becomes stochastic: the ODE’s coefficients Eqs.˜21 and 23 become effectively random due to thermal perturbations in the plasma, and this in turn originates into random initial conditions for the ODE. As a result, the integration of the ODE must be interpreted over an ensemble of particles, and one has to shift from the deterministic view of the cold model, where the bubble radius $r_{b}(\xi)$ was representing the "trajectory of the innermost electron", to a warm interpretation where the bubble radius is now a statistical average $\langle r_{b}(\xi)\rangle$ of an ensemble of randomly sampled trajectories. This stochasticity is driven by the nature of thermal plasma electrons, that even at rest possess non-zero momenta that can be sampled from thermal equilibrium distributions.
@@ 3 Model Characterization and Comparison with Numerical Simulations
In this section, we will characterize the model by solving the blowout radius equation Eq.˜20. Furthermore, we will present comparisons with PIC simulations performed using FBPIC [27], a quasi-3D solver that employs a pseudo-spectral decomposition of the fields in Fourier-Hankel space, enabling tunable azimuthal resolutions, and which also includes standard PIC components for particle evolution like particle pushers [38] and charge deposition schemes [2, 13, 50]. All numerical results presented consider a domain of size $L_{r}=3.0$ and $L_{z}=15.0$, with cell resolutions $\Delta_{z}=\Delta_{r}=0.01$, and computational time step $\Delta t=0.001$. The background plasma is modeled using $64$ particles per cell, and it is initially set at uniform density $n_{0}=10^{16}\;\rm{cm^{-3}}$. In PIC simulations, the background temperature is implemented through a finite variance in the initial particle velocity distribution. We adopt the bi-Gaussian bunch described in Eq.˜1, with $\alpha=84$, $\sigma_{r}=0.1$, and $\sigma_{z}=\sqrt{2}$, in order to satisfy the narrow bunch condition described at the end of Section˜2.2. The choice of parameters grants a complete blowout regime (the normalized charge parameter  [^2] For the bi-Gaussian bunch detailed in Eq. 1, one defines the normalized charge parameter as $\tilde{Q}=\alpha(2\pi)^{3/2}\sigma_{r}^{2}\sigma_{z}$ [39]. , $\tilde{Q}$, which is often used to evaluate linearity regimes in PWFA applications, equals $\sim 18$) and adhere to the parameters chosen in [32, 30].
Figure 5: Parametric scan of the electron sheath width parameters $(\epsilon,\Delta_{l})$, shown through color maps of the relative error metric $C(\epsilon,\Delta_{l})$, as defined in Eq.˜28. Each panel corresponds to a different background temperature: ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=0.0,1.5$ and $3.0\;\rm{keV}$ (panels from top to bottom). Dashed white lines represent $10\%$ error curves.
In Fig.˜1, we show a snapshot of the system under consideration coming from a PIC simulation. The upper half of the figure corresponds to a cold plasma setup, while the lower half shows the warm plasma case. A clear difference emerges between the two: in the presence of a finite plasma temperature, the plasma bubble becomes longitudinally smaller. This reduction in bubble size is expected from linear plasma theory, where finite thermal effects modify the plasma wave frequency through the Bohm-Gross dispersion relation [4, 6], leading to a reduced plasma wavelength [46, 47]. In the blowout regime considered here, however, linear arguments do not longer apply, and hence the longitudinal contraction needs to be quantitatively inspected. This effect can be characterized by looking at the spatial extent of the interface $r_{b}(\xi)$, the boundary separating the interior of the bubble from the surrounding plasma sheath, where electrons accumulation occurs. In warm plasmas, this sheath is broader and less sharply defined compared to the cold case [21]. This is a direct consequence of the statistical spread in electron momenta due to thermal fluctuations.
Figure 6: Numerical solutions of Eq.˜20 compared with PIC results for two initial background temperatures ( $k_{B}T_{i}=0.00\;\rm{keV}$ and $1.00\;\rm{keV}$). In the cold case (top half), the trajectory $r_{b}(\xi)$ is deterministic; in the warm case (bottom half), we show the ensemble average $\langle r_{b}(\xi)\rangle$. In both cases, the sheath width $\Delta$ is selected according to the parameter scan of Fig.˜5.
@@
In order to investigate how the theoretical model exposed in Section˜2 captures this behavior, we numerically solve the ODE in Eq.˜20 for an ensemble of $500$ stochastic realizations of the initial quantities $E^{(0)}$ and $p_{z}^{(0)}$. These quantities are sampled according to a thermal distribution, i.e., a normal distribution $\mathcal{N}(\nu,\sigma^{2})$ with mean $\nu$ and variance $\sigma^{2}$, using the relations:
@@
This sampling reflects first on the form of the coefficients of the ODE Eqs.˜21 and 23. Furthermore, each sampled state yields a different initial condition for the trajectory solver. From Eq.˜11 one gets:
@@
using the same parameterization as in Lu’s model [32, 30], in order to evaluate temperature effects independently of possible changes due to a different choice for the source term $S$. We remark that in Eq.˜26, $\epsilon$ represents the $r_{b}$-dependent part of the electron sheath, while $\Delta_{l}$ is its constant part, according to Lu’s own definition. In panel (a) of Fig.˜2, we show the solution of Eq.˜20 for an initial plasma temperature of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=1\rm{keV}$. The stochasticity of the initial conditions gives rise to an ensemble of trajectories (shown in gray), that represents in this model the sheath thermal spreading also observed in [21], and from which we define the representative bubble trajectory as their statistical average, $\langle r_{b}(\xi)\rangle$ (shown in orange). In panel (b), we compare results coming from a cold realization of the model with the warm ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=1\rm{keV}$ case. One observes that the statistical average $\langle r_{b}(\xi)\rangle$ (dashed orange line) shows only modest deviations from the cold bubble trajectory (shown in blue), indicating that the averaging process alone is not sufficient to fully capture thermal effects, at least at the temperature values considered here. Nevertheless, the stochastic fluctuations that arise from the ODE are not negligible. In panel (c) we quantify these fluctuations through
@@
evaluated at $\xi=11.0$, and computed for different values of the initial background temperature. As anticipated from Section˜3, the fluctuations exhibit a $\sim({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i})^{1/2}$ scaling, reminiscent of the behavior found in the expression for the Debye length. We note however that this analogy is only formal, as Debye shielding appears only in the presence of thermal equilibration, and such effect does not appear in this context.
@@
A similar analysis is presented in Fig.˜3 for the first derivative $r_{b}^{\prime}$, which can be related via Eq.˜11 to momentum variables, and thus its ensemble fluctuations are naturally connected to thermal spread. Panels (a) and (b) lead to the same conclusions as in the previous figures: with the parametrization of Eq.˜26, the stochastic averages deviate only slightly from the cold solution. On the other hand the fluctuations remain significant at both the head and tail of the bubble, and when plotting the metric $\Delta r_{b}^{\prime}(\xi)$, defined analogously to Eq.˜27, we observe a trend that echoes the temperature profile reported in mono-dimensional studies [11], with a rise at the bubble front and an even more pronounced peak at its tail.
@@
The presence of thermal fluctuations implies a non-trivial impact on the structure of the electron sheath that defines the edge of the bubble and hence, on the effective source term itself. Such effect is clearly visible in Fig.˜4, where we show results from PIC simulations on the radial cross section of the source term $S$ at a fixed value of the co-moving coordinate, $\xi_{0}=7.0$, corresponding to the longitudinal center of the bi-Gaussian bunch. As the background temperature increases, the region of electron accumulation at the boundary of the bubble broadens and its peak amplitude decreases. Indeed, this behavior suggests that an accurate description of the warm plasma regime requires not only the inclusion of stochastic initial conditions in the coefficients $E^{(0)}$ and $p_{z}^{(0)}$, but also in a temperature dependent parametrization of the source term that captures the underlying statistical spread induced by thermal effects. A derivation of such parametrization from first principles is indeed challenging and worthy of future investigations. As a preliminary step in this direction, we have performed a parameter-space scan over the terms $(\epsilon,\Delta_{l})$ defined in Eq.˜26 to identify, a posteriori, the most suitable representation of the electron sheath width $\Delta$, and hence of the source term $S$. To this end, we have introduced a metric that quantifies the discrepancy between the bubble trajectory $r_{b}^{(sim)}$ obtained from PIC simulations and a theoretical realization $\langle r_{b}\rangle(\epsilon,\Delta_{l})$ of the model Eq.˜20, corresponding to a given set of parameters:
@@
The results of this parametric scan are shown in Fig.˜5 for three values of background temperature, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=0.0,1.5,3.0\;\rm{keV}$. Comparisons of the $10\%$ error curves (depicted as dashed white lines in the figure) show that an increase in temperature implies a change in the optimal values of both $\Delta_{l}$ and $\epsilon$. The single-sheath parametrization of the source term adopted by [32, 30] and in this work implies that the contribution of the source term $S$ to the pseudo-potential $\psi$ can be effectively divided into three distinct regions: an inner cavity, a narrow electron sheath, and an outer linear-response region where plasma perturbation is minimal. One then sees that as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}$ increases, the $10\%$ error curves marking the optimal parameters region shifts downward. This shows that the optimal $\Delta_{l}$, expected to be on the order of the cold plasma wavelength $k_{p}^{-1}$ in the cold limit (equal to $1$ in dimensionless coordinates [32, 30]), decreases. This trend can be interpreted as a consequence of the overall compression of characteristic length scales in warm plasmas compared to cold ones [4, 6]. Additionally, the parameter $\Delta_{s}=\epsilon r_{b}$ sets the width of the sheath region, where the source term rises sharply. As shown in Fig.˜4, the height of this region diminishes with increasing temperature; charge conservation then requires a compensating increase in its width. In Fig.˜5, this is reflected into a "thermal shrinkage" of the $10\%$ error curves at low $\epsilon$, effectively excluding these values from the optimal region.
Figure 7: Comparison between PIC simulations (in orange) and the thermal model (in blue). The top panel shows the cold normalized quantity $\ell_{\parallel}/\ell_{\parallel}^{cold}$ versus the initial background temperature, while the bottom panel highlights the $\sim(k_{B}T_{i})^{1/2}$ scaling (represented in dashed black as a reference) of the quantity $1-\ell_{\parallel}/\ell_{\parallel}^{cold}$
In Fig.˜6 we report on the combined effect of the new source term parametrization and the stochastic averaging procedure introduced earlier, showing that they yield a reasonable reconstruction of the interface $r_{b}$ in the warm-plasma case (bottom half of the figure). In particular, the model successfully captures the reduction in the longitudinal extent of the bubble relative to the cold-plasma case (top half), a consequence of thermal effects. To deepen the analysis, one could quantify how the bubble size varies with temperature to reveal possible scaling laws. Bubble size can be parametrized via the longitudinal ($\ell_{||}$) and transversal ($\ell_{\perp}$) length scales (see Fig.˜1) [48]. Starting from the maximum radial elongation $r_{m}$ and the corresponding co-moving coordinate $\xi_{m}$, the two observables are defined as $\ell_{\parallel}=\xi_{\parallel}-\xi_{m}$ and $\ell_{\bot}=r_{m}=r_{b}\left(\xi=\xi_{m}\right)$, where $\xi_{\parallel}$ is the point on the longitudinal axis where the bubble ends. Moreover, since transversal observables are less sensitive to the temperature [48], we decided to focus on the characterization of $\ell_{||}$ as a function of $k_{B}T_{i}$.Results of the analysis are presented in Fig.˜7: in the top panel it is shown that, by carefully selecting the sheath-width parameters within the $10\%$ uncertainty band shown in Fig.˜5, it is possible to reproduce the PIC results for $\ell_{\parallel}$ using the model, for all the ranges of temperature considered. Furthermore, the bottom panel shows that both the model and the simulations produce a relative contraction following a scaling law of the form
$1-\ell_{\parallel}/\ell_{\parallel}^{cold}\sim(k_{B}T_{i})^{1/2}\;,$ (29)
which differs from the contraction one would infer from linear plasma theory [4, 6]. This square-root dependence is closely related to the thermal broadening of the electron sheath already observed in Fig.˜2.
We note that the magnitude of the thermal effects on bubble size may increase for more linear regimes (corresponding to smaller values of $\tilde{Q}$). However, in such cases, the assumptions underlying the model presented in this work, namely the requirement of a complete blowout, would no longer hold. We therefore leave the assessment of this thermal model’s performance in a more linear wakefield regime to future studies.
Lastly, in Fig.˜8 we report the values of the on-axis longitudinal electric fields, for both a cold (top panel) and a warm (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=1\;\rm{keV}$, bottom panel) background plasma. PIC results are confronted against solutions of the theoretical model presented in this article (that can easily be computed from $r_{b}^{\prime}$ [32, 30]). It is observed that the model is able to reproduce accurately the field inside the bubble, with noticeable deviations only at its tail (where the interface $r_{b}(\xi)$ becomes non derivable). Furthermore, the structure of the longitudinal field is not significantly altered by the warm background plasma, as already expected from previous studies [21]. A more quantitative analysis of the importance of thermal effects on electromagnetic fields can be found in [48].